# Locally recoverable algebro-geometric codes from projective bundles

### v2

@@ 3.4. Minimum distance of ${\mathcal{C}}$ for small locality
Let $b\geq 3$ and $r=1$, $2$, or $3$. Then the code $\mathcal{C}$ is an optimal LRC with parameters $[n,k,d;r]_{q}$, where
@@ 3.5. Minimum distance of ${\mathcal{C}}$ for localities $r\geq 4$
The inequality (3.3) in the proof of Lemma 3.10 does not hold when $r\geq 4$. Instead, we have
$(b-2)(r+1)+(r-1)\leq b(r-1)+2s\quad\text{for }r\geq 4,$
and the right-hand side increases as $s$ increases. We thus lose control of the lower bound on the minimum distance of ${\mathcal{C}}$, opening up the possible existence of nonoptimal codes ${\mathcal{C}}$ when $r\geq 4$. We exhibit one such code in detail in Section 4. Nevertheless, we can still prove that, given a large alphabet, most choices of points for ${\mathcal{C}}$ yield optimal LRCs.
Suppose that $\mathbf{c}\in{\mathcal{C}}$ is a nonzero code word of weight $\leq r+2$, with $s$ zero-fibers; permuting indices if necessary, we may assume that the (disjoint) union of these zero fibers is $A_{1}\amalg\cdots\amalg A_{s}$. In particular, for $1\leq i\leq s$, the single-variable polynomial $f(x_{i},y)=\sum_{\ell=0}^{r-1}a_{\ell}(x_{i})y^{\ell}\in{\mathbb{F}}_{q}[y]$ of degree $\leq r-1$ vanishes at $y_{i,1},\dots,y_{i,r+1}$. Thus $a_{\ell}(x)$ vanishes at $x_{1},\dots,x_{s}$ for all $10\leq\ell\leq r-$ 1. Hence
@@
$(b-s)(r+1)-(r+2)=(b-s-1)r+(b-s-2)$
points in the $b-s$ fibers over $x_{s+1},\dots,x_{b}$. However, $g(x,y)$ has only $(b-s-1)r$ coefficients as a polynomial in ${\mathbb{F}}_{q}[x,y]$. On the other hand, by Corollary 3.12, we know that $s\leq b-3$, so that $b-s-2\geq 1$. Thus, if we think of the coefficients of $g(x,y)$ as $(b-s-1)r$ unknowns satisfying $(b-s-1)r+(b-s-2)$ linear relations of the form $g(x_{i},y_{i,j})=0$, a code word $\mathbf{c}\in{\mathcal{C}}$ of weight $\leq r+2$ solves a seemingly overconstrained linear system of equations. We should expect that, for most choices of points ${\mathcal{P}}$ with distinct $x$- and $y$-coordinates, it is not possible to solve this system of linear equations. However, there may be polynomials of the form (3.4) with $s$ zero-fibers in multiple ways, and we must take into account polynomials whose zero-fibers are in arbitrary positions (not just over $x_{1},\dots,x_{s}$). We can construct these by varying the roots of the largest factor of $f(x,y)$ that depends only on $x$ and changing the $g(x,y)$, while the collection of points ${\mathcal{P}}$ remains fixed. This gives new opportunities for $(b-s-1)r+(b-s-2)$ of the remaining $(b-s)(r+1)$ points to interpolate the polynomial $g(x,y)$. More precisely, there are
@@
choices of subsets in ${\mathcal{P}}$ consisting of $s$ fibers and $(b-s-1)r+(b-s-2)$ points in the remaining $(b-s)$ fibers. We must now estimate the probability that, given one such subset, there exists a nonzero polynomial $f(x,y)\in V$ vanishing along the subset, giving a code word of weight $\leq r+2$. This leads to an estimate of the expected number of code words in ${\mathcal{C}}$ of weight $\leq r+2$. To help fix ideas, we first illustrate this estimate in Example 3.14.
@@
is the unique zero-fiber of $f(x,y)$, so $s=b-3=1$ in this case. Then the polynomial $g(x,y)$ takes the form
@@
If $g(x,y)$ were to pass through the $(b-s-1)r+(b-s-2)=9$ points
@@
#$\frac{\#X({\mathbb{F}}_{q})}{\#U({\mathbb{F}}_{q})}\leq\frac{Aq^{9}}{q^{10}+O(q^{9})}\underset{q\to\infty}{\xrightarrow{\hskip 20.00003pt}}\frac{A}{q}.$
@@
This estimate is coarse. For example, it ignores the remaining eight minors. Nevertheless, it approaches 00 as $q\to\infty$. Thus, as $q\to\infty$, the code ${\mathcal{C}}$ will contain no words of weight $\leq 6$ with probability $1$.
@@
Our discussion so far shows that if $f(x,y)\in V$ gives rise to a nonzero code word $\mathbf{c}\in{\mathcal{C}}$ of weight $\leq r+2$, then $f(x,y)$ vanishes along $s\leq b-3$ zero-fibers, and along at least $(b-s-1)r+(b-s-2)$ points in the remaining $b-s$ fibers. The expression (3.5) quantifies the number of subsets of ${\mathcal{P}}$ along which such an $f(x,y)$ might vanish. Let ${\mathcal{P}}^{\prime}$ be one of these subsets, partitioned as ${\mathcal{P}}^{\prime}={\mathcal{P}}_{1}\amalg{\mathcal{P}}_{2}$, where ${\mathcal{P}}_{1}=A_{i_{1}}\amalg\cdots\amalg A_{i_{s}}$ are the $s$ zero-fibers above $x_{i_{1}},\dots,x_{i_{s}}$ and ${\mathcal{P}}_{2}$ contains $(b-s-1)r+(b-s-2)$ points in the remaining $b-s$ fibers. Note that $f(x,y)$ is allowed to vanish at points of ${\mathcal{P}}\setminus{\mathcal{P}}^{\prime}$, as long is it does not vanish along zero-fibers not already contained in ${\mathcal{P}}_{1}$. Then
@@
for some $g_{\ell}(x)\in{\mathbb{F}}_{q}[x]$ of degree $\leq b-s-2$ and $10\leq\ell\leq r-$ 1. The number of distinct $x$-coordinates among the points in ${\mathcal{P}}_{2}$ is $u\coloneqq b-s$, by definition of $s$, and the number of distinct $y$-coordinates is $v\coloneqq\#{\mathcal{P}}_{2}=(b-s-1)r+(b-s-2)$. Just as in (3.7), the condition that $g(P)=0$ for all $P\in{\mathcal{P}}_{2}$ can be written as a matrix equation
@@
matrix in $u+v=(b-s-1)(r+2)$ variables, and $\mathbf{a}$ encodes the coefficients of $g(x,y)$. If $f(x,y)$ is not the zero polynomial, the matrix $M$ must have a nontrivial kernel, which means that all its $\binom{(b-s-1)r+(b-s-2)}{(b-s-2)}$ maximal minors must vanish. Each minor is a homogeneous polynomial in $N\coloneqq u+v-(b-s-2)$ variables, because each row removed from $M$ to obtain a maximal minor reduces the total number of variables by one. Thus, each maximal minor defines a hypersurface $X\subset{\mathbb{P}}^{N-1}_{{\mathbb{F}}_{q}}$. The Lang–Weil estimate (Theorem 2.1) implies that
#$\#X({\mathbb{F}}_{q})\leq Aq^{N-2},$
@@
#$\#{\mathbb{P}}^{N-1}({\mathbb{F}}_{q})=\frac{q^{N}-1}{q-1}=q^{N-1}+q^{N-2}+\cdots+q+1.$
@@
#$\#U({\mathbb{F}}_{q})\geq\frac{q^{N}-1}{q-1}-B\cdot\frac{q^{N-1}-1}{q-1}=q^{N-1}+O(q^{N-2})$
@@
(3.9) #$\frac{\#X({\mathbb{F}}_{q})}{\#U({\mathbb{F}}_{q})}\leq\frac{Aq^{N-2}}{q^{N-1}+O(q^{N-2})}\underset{q\to\infty}{\xrightarrow{\hskip 20.00003pt}}\frac{A}{q}.$
@@
#$\frac{\#X({\mathbb{F}}_{q})}{\#U({\mathbb{F}}_{q})}\sum_{s=0}^{b-3}\binom{b}{s}\binom{(b-s)(r+1)}{(b-s-1)r+(b-s-2)}.$
Now (3.9) guarantees that this quantity approaches 00 as $q\to\infty$, which finishes the proof. ∎
@@
    It would be interesting to make the proof of Theorem 3.15 effective. That is, for a given pair of thresholds $10<\gamma_{1},\gamma_{2}<$ 1, obtain an explicit estimate of how large $q_{0}=q_{0}(\gamma_{1},\gamma_{2})$ must be so that a proportion $\geq\gamma_{1}$ of the choices for ${\mathcal{P}}$ yields an optimal LRC with probability $\geq\gamma_{2}$.
@@ 3.6. A detour through matroids
The rank of a circuit $L$ in a matroid $\mathsf{M}$ is #$\#L-1$. Given $\varphi\colon\mathsf{M}\to\mathsf{N}$ an isomorphism of matroids, then $L$ is a circuit in $\mathsf{M}$ if and only if $\varphi(L)$ is a circuit in $\mathsf{N}$.
@@
Set $d_{{\mathcal{C}}}$ and $d_{{\mathcal{C}}^{\prime}}$ the minimum distances of ${\mathcal{C}}$ and ${\mathcal{C}}^{\prime}$, respectively. By Lemma 3.19 there are circuits $L^{\bot}$ in $\mathsf{M}_{{\mathcal{C}}^{\bot}}$ and $L^{\prime\bot}$ in $\mathsf{M}_{{\mathcal{C}}^{\prime\bot}}$ such that $d_{{\mathcal{C}}}=\#L^{\bot}$ and $d_{{\mathcal{C}}^{\prime}}=\#L^{\prime\bot}$. The isomorphism of matroids $\mathsf{M}_{{\mathcal{C}}}\cong\mathsf{M}_{{\mathcal{C}}^{\prime}}$ induces an isomorphism between the dual matroids $(\mathsf{M}_{{\mathcal{C}}})^{\bot}\cong(\mathsf{M}_{{\mathcal{C}}^{\prime}})^{\bot}$. This induces an isomorphism $\varphi:\mathsf{M}_{{\mathcal{C}}^{\bot}}\cong(\mathsf{M}_{{\mathcal{C}}})^{\bot}\cong(\mathsf{M}_{{\mathcal{C}}^{\prime}})^{\bot}\cong\mathsf{M}_{{\mathcal{C}}^{\prime\bot}}$ where the first and third isomorphisms occur by [JurriusPellikaan13]*p. 269. Thus $\varphi^{-1}(L^{\prime\bot})$ is a circuit in $\mathsf{M}_{{\mathcal{C}}^{\bot}}$ and $\varphi(L^{\bot})$ is a circuit in $\mathsf{M}_{{\mathcal{C}}^{\prime\bot}}$ by Remark 3.17. Since $L^{\bot}$ and $L^{\prime\bot}$ have the smallest cardinality among the circuits in $\mathsf{M}_{{\mathcal{C}}^{\bot}}$ and $\mathsf{M}_{{\mathcal{C}}^{\prime\bot}}$ respectively, then #$\#L^{\bot}\leq\#\varphi^{-1}(L^{\prime\bot})=\#L^{\prime\bot}$ and #$\#L^{\prime\bot}\leq\#\varphi(L^{\bot})=\#L^{\bot}$. Thus $d_{{\mathcal{C}}}=\#L^{\bot}=\#L^{\prime\bot}=d_{{\mathcal{C}}^{\prime}}$. ∎
@@
The rank of $M$ and $M^{\prime}$ is $b-1$, so each of their kernels is one dimensional. Set

### v3

@@ 3.4. Minimum distance of $\mathcal{C}$} for small locality
Let $b\geq 3$ and $r=1$, 22, or $3$. Then the code $\mathcal{C}$ is an optimal LRC with parameters $n,k,d;r]_{q}$, where
@@ 3.5. Minimum distance of $\mathcal{C}$} for localities $r\geq 4$
The inequality (3.3) in the proof of Lemma 3.10 does not hold when $r\geq 4$. We thus lose control of the lower bound on the minimum distance of $\mathcal{C}$}, opening up the possible existence of nonoptimal codes $\mathcal{C}$} when $r\geq 4$. We exhibit one such code in detail in Section 4. Nevertheless, we can still prove that, given a large alphabet, most choices of points for $\mathcal{C}$} yield optimal LRCs.
Suppose that $\mathbf{c}\in{\mathcal{C}}$ is a nonzero code word of weight $\leq r+2$, with $s$ zero-fibers; permuting indices if necessary, we may assume that the (disjoint) union of these zero fibers is $A_{1}\amalg\cdots\amalg A_{s}$. In particular, for $1\leq i\leq s$, the single-variable polynomial $f(x_{i},y)=\sum_{\ell=0}^{r-1}a_{\ell}(x_{i})y^{\ell}\in{\mathbb{F}}_{q}[y]$ of degree $\leq r-1$ vanishes at $y_{i,1},\dots,y_{i,r+1}$. Thus $a_{\ell}(x)$ vanishes at $x_{1},\dots,x_{s}$ for all $0\leq\ell\leq r-1$. Hence
@@
(bs)(r+1)(r+2)=(bs1)r+(bs2)(b-s)(r+1)-(r+2)=(b-s-1)r+(b-s-2)
points in the bsb-s fibers over $x_{s+1},\dots,x_{b}$. However, $g(x,y)$ has only (bs1)r(b-s-1)r coefficients as a polynomial in ${\mathbb{F}}_{q}[x,y]$. On the other hand, by Corollary 3.12, we know that $s\leq b-3$, so that $b-s-2\geq 1$. Thus, if we think of the coefficients of $g(x,y)$ as (bs1)r(b-s-1)r unknowns satisfying (bs1)r+(bs2)(b-s-1)r+(b-s-2) linear relations of the form $g(x_{i},y_{i,j})=0$, a code word $\mathbf{c}\in{\mathcal{C}}$ of weight $\leq r+2$ solves a seemingly overconstrained linear system of equations. We should expect that, for most choices of points $\mathcal{P}$} with distinct $x$- and $y$-coordinates, it is not possible to solve this system of linear equations. However, there may be polynomials of the form (3.4) with $s$ zero-fibers in multiple ways, and we must take into account polynomials whose zero-fibers are in arbitrary positions (not just over $x_{1},\dots,x_{s}$). We can construct these by varying the roots of the largest factor of $f(x,y)$ that depends only on $x$ and changing the $g(x,y)$, while the collection of points $\mathcal{P}$} remains fixed. This gives new opportunities for (bs1)r+(bs2)(b-s-1)r+(b-s-2) of the remaining (bs)(r+1)(b-s)(r+1) points to interpolate the polynomial $g(x,y)$. More precisely, there are
@@
choices of subsets in $\mathcal{P}$} consisting of $s$ fibers and (bs1)r+(bs2)(b-s-1)r+(b-s-2) points in the remaining (bs)(b-s) fibers. We must now estimate the probability that, given one such subset, there exists a nonzero polynomial $f(x,y)\in V$ vanishing along the subset, giving a code word of weight $\leq r+2$. This leads to an estimate of the expected number of code words in $\mathcal{C}$} of weight $\leq r+2$. To help fix ideas, we first illustrate this estimate in Example 3.14.
@@
is the unique zero-fiber of $f(x,y)$, so s=b3=1s=b-3=1 in this case. Then the polynomial $g(x,y)$ takes the form
@@
If $g(x,y)$ were to pass through the (bs1)r+(bs2)=9(b-s-1)r+(b-s-2)=9 points
@@
$\frac{\#X({\mathbb{F}}_{q})}{\#U({\mathbb{F}}_{q})}\leq\frac{Aq^{9}}{q^{10}+O(q^{9})}\underset{q\to\infty}{\xrightarrow{\hskip 16.38895pt}}\frac{A}{q}.$
@@
This estimate is coarse. For example, it ignores the remaining eight minors. Nevertheless, it approaches 0 as $q\to\infty$. Thus, as $q\to\infty$, the code $\mathcal{C}$} will contain no words of weight $\leq 6$ with probability $1$.
@@
Our discussion so far shows that if $f(x,y)\in V$ gives rise to a nonzero code word $\mathbf{c}\in{\mathcal{C}}$ of weight $\leq r+2$, then $f(x,y)$ vanishes along $s\leq b-3$ zero-fibers, and along at least (bs1)r+(bs2)(b-s-1)r+(b-s-2) points in the remaining bsb-s fibers. The expression (3.5) quantifies the number of subsets of $\mathcal{P}$} along which such an $f(x,y)$ might vanish. Let $\mathcal{P}$$^{\prime}$ be one of these subsets, partitioned as ${\mathcal{P}}^{\prime}={\mathcal{P}}_{1}\amalg{\mathcal{P}}_{2}$, where ${\mathcal{P}}_{1}=A_{i_{1}}\amalg\cdots\amalg A_{i_{s}}$ are the $s$ zero-fibers above $x_{i_{1}},\dots,x_{i_{s}}$ and ${\mathcal{P}}_{2}$ contains (bs1)r+(bs2)(b-s-1)r+(b-s-2) points in the remaining bsb-s fibers. Note that $f(x,y)$ is allowed to vanish at points of ${\mathcal{P}}\setminus{\mathcal{P}}^{\prime}$, as long is it does not vanish along zero-fibers not already contained in ${\mathcal{P}}_{1}$. Then
@@
for some $g_{\ell}(x)\in{\mathbb{F}}_{q}[x]$ of degree $\leq b-s-2$ and $0\leq\ell\leq r-1$. The number of distinct $x$-coordinates among the points in ${\mathcal{P}}_{2}$ is $u\coloneqq b-s$, by definition of $s$, and the number of distinct $y$-coordinates is $v\coloneqq\#{\mathcal{P}}_{2}=(b-s-1)r+(b-s-2)$. Just as in (3.7), the condition that $g(P)=0$ for all $P\in{\mathcal{P}}_{2}$ can be written as a matrix equation
@@
matrix in u+v=(bs1)(r+2)u+v=(b-s-1)(r+2) variables, and $\mathbf{a}$ encodes the coefficients of $g(x,y)$. If $f(x,y)$ is not the zero polynomial, the matrix $M$ must have a nontrivial kernel, which means that all its $\binom{(b-s-1)r+(b-s-2)}{(b-s-2)}$ maximal minors must vanish. Each minor is a homogeneous polynomial in $N\coloneqq u+v-(b-s-2)$ variables, because each row removed from $M$ to obtain a maximal minor reduces the total number of variables by one. Thus, each maximal minor defines a hypersurface $X\subset{\mathbb{P}}^{N-1}_{{\mathbb{F}}_{q}}$. The Lang–Weil estimate (Theorem 2.1) implies that
$X({\mathbb{F}}_{q})\leq Aq^{N-2},$
@@
${\mathbb{P}}^{N-1}({\mathbb{F}}_{q})=\frac{q^{N}-1}{q-1}=q^{N-1}+q^{N-2}+\cdots+q+1.$
@@
$U({\mathbb{F}}_{q})\geq\frac{q^{N}-1}{q-1}-B\cdot\frac{q^{N-1}-1}{q-1}=q^{N-1}+O(q^{N-2})$
@@
(3.9) $\frac{\#X({\mathbb{F}}_{q})}{\#U({\mathbb{F}}_{q})}\leq\frac{Aq^{N-2}}{q^{N-1}+O(q^{N-2})}\underset{q\to\infty}{\xrightarrow{\hskip 16.38895pt}}\frac{A}{q}.$
@@
$\frac{\#X({\mathbb{F}}_{q})}{\#U({\mathbb{F}}_{q})}\sum_{s=0}^{b-3}\binom{b}{s}\binom{(b-s)(r+1)}{(b-s-1)r+(b-s-2)}$.
Now (3.9) guarantees that this quantity approaches 0 as $q\to\infty$, which finishes the proof. ∎
@@
    It would be interesting to make the proof of Theorem 3.15 effective. That is, for a given pair of thresholds $0<\gamma_{1},\gamma_{2}<1$, obtain an explicit estimate of how large $q_{0}=q_{0}(\gamma_{1},\gamma_{2})$ must be so that a proportion $\geq\gamma_{1}$ of the choices for $\mathcal{P}$} yields an optimal LRC with probability $\geq\gamma_{2}$.
@@ 3.6. A detour through matroids
The rank of a circuit $L$ in a matroid $\mathsf{M}$ is #L1\#L-1. Given $\varphi\colon\mathsf{M}\to\mathsf{N}$ an isomorphism of matroids, then $L$ is a circuit in $\mathsf{M}$ if and only if $\varphi(L)$ is a circuit in $\mathsf{N}$.
@@
Set $d_{{\mathcal{C}}}$ and $d_{{\mathcal{C}}^{\prime}}$ the minimum distances of $\mathcal{C}$} and $\mathcal{C}$$^{\prime}$, respectively. By Lemma 3.19 there are circuits $L^{\bot}$ in $\mathsf{M}_{{\mathcal{C}}^{\bot}}$ and $L^{\prime\bot}$ in $\mathsf{M}_{{\mathcal{C}}^{\prime\bot}}$ such that $d_{{\mathcal{C}}}=\#L^{\bot}$ and $d_{{\mathcal{C}}^{\prime}}=\#L^{\prime\bot}$. The isomorphism of matroids $\mathsf{M}_{{\mathcal{C}}}\cong\mathsf{M}_{{\mathcal{C}}^{\prime}}$ induces an isomorphism between the dual matroids $(\mathsf{M}_{{\mathcal{C}}})^{\bot}\cong(\mathsf{M}_{{\mathcal{C}}^{\prime}})^{\bot}$. This induces an isomorphism $\varphi:\mathsf{M}_{{\mathcal{C}}^{\bot}}\cong(\mathsf{M}_{{\mathcal{C}}})^{\bot}\cong(\mathsf{M}_{{\mathcal{C}}^{\prime}})^{\bot}\cong\mathsf{M}_{{\mathcal{C}}^{\prime\bot}}$ where the first and third isomorphisms occur by [JurriusPellikaan13]*p. 269. Thus $\varphi^{-1}(L^{\prime\bot})$ is a circuit in $\mathsf{M}_{{\mathcal{C}}^{\bot}}$ and $\varphi(L^{\bot})$ is a circuit in $\mathsf{M}_{{\mathcal{C}}^{\prime\bot}}$ by Remark 3.17. Since $L^{\bot}$ and $L^{\prime\bot}$ have the smallest cardinality among the circuits in $\mathsf{M}_{{\mathcal{C}}^{\bot}}$ and $\mathsf{M}_{{\mathcal{C}}^{\prime\bot}}$ respectively, then $L^{\bot}\leq\#\varphi^{-1}(L^{\prime\bot})=\#L^{\prime\bot}$ and $L^{\prime\bot}\leq\#\varphi(L^{\bot})=\#L^{\bot}$. Thus $d_{{\mathcal{C}}}=\#L^{\bot}=\#L^{\prime\bot}=d_{{\mathcal{C}}^{\prime}}$. ∎
@@
The rank of $M$ and $M^{\prime}$ is b1b-1, so each of their kernels is one dimensional. Set